OpenStudy (anonymous):

cos(2x)=cos(x) Hint: start by simplifying with an identity.

OpenStudy (yb1996):

cos(2x) = cos(x) 2cos^2(x) -1 = cos(x) 2cos^2(x) - cos(x) - 1 = 0 2cos^2(x) - 2cosx + cosx - 1 = 0 2cos(x)[cos(x) - 1] + 1[cos(x) - 1] = 0 [cos(x) - 1][2cos(x) + 1] = 0 cosx = 1 or -1/2 x = 0, 2pi , 2pi/3, 4pi/3,